Suppose that X = (Xi, X2, , X.) and Y-X,,Y2, , Ym) are random samples from continuous distributions F and G, respectively.Wilcoxons two-sample test statistic W - W(X, Y) is defined to beRi where R, is the rank of Y in the combined sample 1. Let T Σǐn i Zi where Zi,Z2, ,Zm are numbers sampled at random without replacement from the set {1,2,..., N} Show that E(Z) = (N + 1)/2 and hence E(T) m(N + 1)/2 Show that 3N2 +5N+2 E(Z) = and hence 12 ー1 Deduce that under the null hypothesis that F = G, the expectation and variance of Wilcoxons two-sample test statistic are m(n+m+1)/2 and nm(n+m+1)/12, respectively
Answer #1

, jメj= 1(1) 11N(N +1) N+1 1N(N +1) (2N+1 6 (N +1)(2N +1) 6 2 j判 N(N 1) 12(N-1) 12 12 N2-1

m(N2 1) m(m-1)(N + 1) 12 2- 12 If Zis are rank of Ys in the combined sample then RZ and under Ho F = G then we select Zis, i = 1(1)m randomly from {1.2, ,y without replacement. Here, Ņ = n+m. then put m+n instead of N in E(T) and Var(T), we get expectation m(n+m+1)/2, variance-mn(mn+1)/12.